# A Neural Rewriting System to Solve Algorithmic Problems

### v1

@@ Abstract
Modern neural network architectures still struggle to learn algorithmic procedures that require to systematically apply compositional rules to solve out-of-distribution problem instances. In this work, we propose an original approach to learn algorithmic tasks inspired by rewriting systems, a classic framework in symbolic artificial intelligence. We show that a rewriting system can be implemented as a neural architecture composed by specialized modules: the Selector identifies the target sub-expression to process, the Solver simplifies the sub-expression by computing the corresponding result, and the Combiner produces a new version of the original expression by replacing the sub-expression with the solution provided. We evaluate our model on three types of algorithmic tasks that require simplifying symbolic formulas involving lists, arithmetic, and algebraic expressions. We test the extrapolation capabilities of the proposed architecture using formulas involving a higher number of operands and nesting levels than those seen during training, and we benchmark its performance against the Neural Data Router, a recent model specialized for systematic generalization, and a state-of-the-art large language model (GPT-4) probed with advanced prompting strategies.
@@ 1 Introduction
Whether neural networks can learn to perform systematic reasoning is a long-standing problem in artificial intelligence and cognitive science Fodor and Pylyshyn (1988), and the recent success of deep learning in solving challenging reasoning tasks has reinvigorated this debate Hupkes et al. (2020). However, despite some promising achievements Lake and Baroni (2023), there is general consensus that even state-of-the-art transformer architectures and large language models still lack systematic and compositional generalization skills Csordás et al. (2022a); Press et al. (2023); Dziri et al. (2023).
In this work, we investigate this issue by considering a class of reasoning problems that requires to simplify symbolic formulas to a minimal form. Despite its apparent simplicity (these problems could be effectively tackled by automatic symbolic solvers), learning to solve this kind of problems is extremely challenging, as it implies the discovery of algorithmic procedures that generalize to formulas with different nesting depths and number of operands.
Inspired by rewriting systems from the tradition of symbolic artificial intelligence Dershowitz and Jouannaud (1990) and by the evidence that learning algebra retrains our visual system to perceive the hierarchical structure of symbolic expressions Marghetis et al. (2016), we propose a neural architecture that decomposes formula simplification problems into sub-tasks carried out by specialized neural modules.
We compare our approach to the Neural Data Router Csordás et al. (2022b), an end-to-end trained variant of a transformer encoder designed to solve algorithmic problems, and to GPT-4 OpenAI (2023) using self-consistency zero-shot CoT prompting Wang et al. (2023), which improves performance of large language models on reasoning tasks. We show that our approach is more effective than the alternatives, especially when solving deeply nested formulas. We also analyze the causes of failure of our system: we find that accurately selecting the portion to be simplified in very long formulas is the hardest sub-task, shedding light on the limitations of current approaches to deal with length generalization in algorithmic tasks with transformers.
@@ 2 Related works
Systematic Generalization A consistent stream of research investigates the extent to which neural networks can learn to reason in a systematic and compositional way Hupkes et al. (2020); Testolin (2024). Several benchmarks have been proposed to measure this capacity: among the most popular, ListOps Nangia and Bowman (2018) and CTL Liska et al. (2018) can be seen as instances of formula simplification problems, in which the learning system is tasked to produce the evaluation of a formula which typically requires to compute intermediate results, either implicitly or explicitly. Other formula simplification benchmarks include derivation and integration problems Lample and Charton (2019) or polynomial simplification Agarwal et al. (2021).
Modern neural architectures achieve different degrees of systematic compositional generalization Ontañón et al. (2022); Csordás et al. (2022b); Lake and Baroni (2023). Interestingly, the emergence of specialized modules in Mixture of Experts architectures has been recently studied by Mittal et al. (2022), who found that while modular architectures can outperform monolithic ones in both in- and out-of-distribution testing conditions, highly specialized modules do not emerge from standard end-to-end training without ad-hoc learning biases.
Rewriting systems and neural networks The idea of implementing neural architectures inspired by symbolic rewriting systems was initially explored in unsupervised learning settings, where individual network weights represented string tokens to be rewritten Komendantskaya. (2009). Others implemented a rewriting system for algebraic problems using custom feature engineering methods and a feed-forward neural network Cai et al. (2018), while in Chen and Tian (2019) the authors propose a reinforcement learning-based system that can learn a general rewriting mechanism by first choosing a region to simplify and then the appropriate rewriting rule to apply to the chosen region.
LLMs Large Language Models have recently been tested on complex reasoning tasks, often exhibiting impressive problem solving capabilities. These tasks often include ‘symbolic reasoning tasks’, such as coin flip, last letter concatenation Wei et al. (2022) and boolean variable assignment Anil et al. (2022), which can be solved by applying simple algorithms. However, LLMs still struggle with problems that require systematic generalization Press et al. (2023).
@@ 3 Formula simplification problems
Many compositional generalization benchmarks require to evaluate a formula which can be made arbitrarily long by replacing elementary operands with expressions that have an equivalent value. For example, in ListOps each integer operand can be replaced with an operation on a list of integers which evaluates to that operand.
@@
In this class of problems, formulas $f\in F$ are composed of two semantically distinct elements: operators $O$ and arguments [^1] We also sometimes refer to arguments as operands. We prefer to use the term arguments here to avoid confusion with operators $O$. $A$. Arguments can be either atomic elements $E$, such as integers, which are also the final values of any formula, or other formulas. Therefore, in general formulas have a recursive structure (see Figure 1). Furthermore, since in these problems each formula has always one and only one final value, any formula has the generic compact form $f=o(a_{1},...,a_{n})=e$, where $o\in O$, $e\in E$ and $a_{j}\in F\cup E,~{}\forall j\in[1,n]$. We can define leaf formulas $F^{L}\subset F$ as the subset of all formulas with form $f^{L}=o(a_{1},...,a_{n})\;\ \mathrm{s.t.}\;\ a_{k}\in E,~{}\forall k\in[1,n]$.
@@
Considering a formula as a sequence of symbols $f=f_{1},f_{2},...,f_{N}$, we can identify the sub-sequences $f^{L}=f_{j},f_{j+1},...,f_{k},j<k$ which correspond to leaf formulas. For any problem, we can further define the set of rewriting rules $r\in R$ which map leaf formulas to their values: $r:F^{L}\rightarrow E,~{}\forall r\in R$. In these problems, some assumptions are made about rewriting rules in $R$. First, rules are assumed to apply to contiguous sub-sequences of symbols. Second, each rewriting rule is assumed to be contractive, that is to map sub-sequences to shorter ones. Third, the rewriting rules have mutually exclusive applicability scope, i.e. given a leaf formula, one and only one rule can be applied to reduce it. Fourth, several leaf formulas could be present in a formula at once, but the order in which rewriting rules are applied on them is assumed to be irrelevant.
According to the formalization outlined above, any algorithmic task which falls into this class of problems can be solved by iterative rewriting the initial formula applying a finite sequence of valid rewriting rules. In general, if at any point in the iterative process there is no valid rule that can be applied to a leaf formula, the computation ends. More precisely, we can identify three sub-tasks which should be performed to solve these problems: selecting a sequence of symbols which corresponds to a leaf formula; solving the leaf formula applying an appropriate rewriting rule; and replacing the leaf formula with the output of the rewriting rule in the original formula.
The task we are interested in this paper consists in learning from a set of (short) formulas with reductions =$\mathcal{F}=\{(f_{i},e_{i})\}_{i=1}^{d}$, where $e_{i}\in E$ is the reduction of formula $f_{i}$, a rewriting system capable of strong out-of-distribution generalization, i.e. able to correctly reduce formulas that are much more complex of the ones in the training set. In addition, we are interested in understanding how far a structured neural architecture can be successful in completing that task.
@@ 4 Neural Rewriting System
Figure 1: Examples of solution of ListOps, Arithmetic and Algebra formulas. Leaf formulas are highlighted in yellow.
@@
In this section, we describe the proposed Neural Rewriting System, a neural architecture composed of three integrated modules, i.e., Selector, Solver, Combiner, where each module is responsible for the solution of a sub-task required to solve a formula by iterative rewriting. A schematic representation of the system is shown in Figure 2.
4.1 Selector Module
In analogy to what happens in humans when they deploy object-based attention to locate algebraic sub-expressions that can be simplified Marghetis et al. (2016), the Selector is trained to identify the last leaf formula [^2] To enable generalization on formulas with more arguments than seen during training, if the identified leaf formula has more than two arguments, we train the Selector to output the smallest fraction of the leaf formula which can be simplified. This corresponds to the operator and the two first arguments appearing in the leaf formula. occurring in the input formula on which a rewriting rule can be applied. When training the Selector, we assume that it will always receive syntactically correct formulas, i.e. we use teacher forcing. The identified leaf formula is then given in input to the Solver. At inference time, the output of the Selector is also used to dynamically set the filters of a Convolutional Neural Network that is a component of the Combiner. Therefore, this module can also be seen as a HyperNetwork Ha et al. (2017) which sets the weights of the Combiner with discrete values.
Since we represent problem instances as strings of text, we frame the problem as a sequence-to-sequence task and use a variant of the transformer Vaswani et al. (2017) to implement the Selector. We modify the vanilla transformer in two ways. First, given the class of problems we consider in this work, we can assume that the problem the Selector needs to solve has a local nature: the leaf formula to be simplified is always located in a limited region of the input sequence. We therefore mask all entries in the self-attention matrix of the encoder but the ones around the main diagonal (i.e., we make them $-\mathrm{inf}$). The diagonal window is $12k+$ 1 tokens wide, where $k$ is an hyperparameter. Second, to improve generalization to out-of-distribution problem instances we require the Selector to be able to identify a leaf formula independently of the length of the input sequence, provided that it is syntactically correct. Following recent evidence that the ability of transformers to generalize on longer samples can be influenced by the choice of positional encodings Ruoss et al. (2023); Csordás et al. (2021), we use Label-based Positional Encodings Li and McClelland (2022) to improve the capability of the Selector to identify leaf formulas in very long sequences. The positional information of an input sequence of $L$ tokens is thus encoded in the following way: given a sequence of $N$ sinusoidal positional encodings, where $N$ is a large number that represents the maximum expected length of an input, $L$ integers are sampled in the interval $[0,N-1]$ and then sorted. The encodings found in the positions corresponding to the sampled integers are then summed to the embeddings of the tokens in the input sequence.
@@ 4.1 Selector Module
Multi-output generation To further increase the accuracy of the Selector on longer sequences, we employ a multi-output generation strategy. We produce several output sequences sampling from the probability distribution derived from the decoder’s outputs and then we choose the best one taking into account both a measure of confidence of the Selector and a measure of input-output agreement computed by the Combiner.
@@
Specifically, we can see each output of the Selector as a candidate leaf formula $\hat{f}^{L}$. We generate any token $\hat{f}_{i}^{L}$ in an output sequence $\hat{f}^{L}$ by sampling from the probability distribution obtained applying the $\mathrm{softmax}$ function to the logits produced by the final fully-connected layer of the decoder. We do not use any temperature parameter when sampling the output tokens. For any input formula $f$, we repeat the stochastic generation process $M$ times, thus generating a sequence of candidate leaf formulas $\hat{F}^{L}=\langle\hat{f}^{L,1},...,\hat{f}^{L,M}\rangle$. We define the confidence of the Selector on any $\hat{f}^{L,j},1\leq j\leq M$ as the joint probability of sampling its tokens: $c(\hat{f}^{L,j})=\prod_{i=1}^{N}p_{i}^{j}$, where $N$ is the number of tokens in $\hat{f}^{L,j}$, and $p_{i}^{j}$ is the probability to sample token $\hat{f}_{i}^{L}$ in $\hat{f}^{L,j}$. We also define an agreement score $a(\hat{f}^{L,j},f)\in[0,1]$ which gives information about the fraction of $\hat{f}^{L,j}$ which is exactly present in the input formula $f$. This measure is computed by the Combiner and thus it is formally defined in Section 4.2. We then select the final output $f^{L}$ of the Selector as the one with the highest Selector confidence which has an agreement score equal to $1$ — that is, it matches the input sequence exactly. More formally, $f^{L}=\hat{f}^{L,j}\in\hat{F}^{L}\;$ $\mathrm{s.t.}\;\ c(\hat{f}^{L,j})\geq c(\hat{f}^{L,k})~{}\forall j,k\in[1,M]\land a(\hat{f}^{L,j},f)=1$.
@@
Dynamic windowing We implement a dynamic windowing mechanism on longer input sequences that allows us to increase the model’s generalization capacity on complex problem instances. We apply this mechanism on top of the multi-output generation strategy by modifying its behavior for sequences longer than a given threshold $T$. Given an input formula $f$, if $|f|<T$ the computation is executed as described in the previous paragraph. Otherwise, we define a sequence of $M$ copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$ corresponding to the $M$ candidate leaf formulas $\langle\hat{f}^{L,1},...,\hat{f}^{L,M}\rangle$ generated by the Selector. The length of each $f^{(k)}$ will be reduced by applying a window function $w$. Considering any input $f$ as a sequence $f_{1},...,f_{N}$ of $N$ tokens, we define the window function $w(f,k)=f_{k+1},...,f_{N}$ which reduces the length of the input by giving as output its last $k$ tokens. Since the Selector is trained to output the last leaf formula appearing in the input, the window function reduces the input length starting from the first tokens. We divide the sequence of copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$ into 10 groups $F^{(1)},...,F^{(10)}$ of equal size. Intuitively, in each group the length of the input is reduced by a different percentage of tokens. More formally, the window function will be parameterized by $k=\mathrm{floor}(|f^{(i)}|\cdot\frac{j}{10})~{}\forall f^{(i)}\in F^{(j)},~{}\forall j\in[1,10]$. This ensures that the model can observe the whole input sequence and also select a leaf expression in the portion of the sequence where it can identify one with highest confidence.

### v2

@@ Abstract
Modern neural network architectures still struggle to learn algorithmic procedures that require to systematically apply compositional rules to solve out-of-distribution problem instances. In this work, we focus on formula simplification problems, a class of synthetic benchmarks used to study the systematic generalization capabilities of neural architectures. We propose a modular architecture designed to learn a general procedure for solving nested mathematical formulas by only relying on a minimal set of training examples. Inspired by rewriting systems, a classic framework in symbolic artificial intelligence, we include in the architecture three specialized and interacting modules: the Selector, trained to identify solvable sub-expressions; the Solver, mapping sub-expressions to their values; and the Combiner, replacing sub-expressions in the original formula with the solution provided by the Solver. We benchmark our system against the Neural Data Router, a recent model specialized for systematic generalization, and a state-of-the-art large language model (GPT-4) probed with advanced prompting strategies. We demonstrate that our approach achieves a higher degree of out-of-distribution generalization compared to these alternative approaches on three different types of formula simplification problems, and we discuss its limitations by analyzing its failures.
[…]
2672
@@ 1 Introduction
Whether neural networks can learn to perform systematic reasoning is a long-standing question in artificial intelligence and cognitive science [10], and the recent success of deep learning on challenging reasoning tasks has reinvigorated this debate [11]. However, despite some promising achievements [15], there is general consensus that even state-of-the-art transformer architectures and large language models still lack systematic and compositional generalization skills [6, 9, 28].
In this work, we investigate this issue by considering a type of systematic generalization benchmark that requires simplifying symbolic formulas to a minimal form. Specifically, we consider the problem of simplifying nested formulas of operations on lists of integers [22], as well as arithmetic and algebraic expressions. In principle these problems could be tackled by automatic symbolic solvers, but here we consider them as benchmarks to assess the reasoning capabilities of trained neural networks under controlled settings [7, 14].
We formalize formula simplification problems and propose a simple solution method inspired by rewriting systems rooted in the tradition of symbolic artificial intelligence [8]. We implement this solution method as a modular neural architecture called the Neural Rewriting System, which represents a case study of a neural architecture in which the roles and interaction mechanisms between neural modules are clearly defined, allowing greater interpretability of the resulting system. We exploit our framework to further evaluate two neural architectures that have been recently shown able to tackle reasoning tasks: the Neural Data Router [7], a transformer encoder trainable end-to-end to solve algorithmic problems, and GPT-4 [24] probed using self-consistency zero-shot Chain-of-Thought [32], which is an advanced prompting method that improves the performance of large language models on reasoning tasks.
Through extensive experiments and analyses, we show that the proposed modular architecture achieves greater accuracy compared to advanced alternative approaches, especially when solving deeply nested formulas. Furthermore, we show that our approach allows us to understand in detail what parts of a systematic generalization problem could represent the greatest obstacles for neural architectures. Indeed, when analyzing the causes of failure of the Neural Rewriting System, we find that accurately selecting the portion to be simplified in very long formulas is the most challenging task for the architecture, shedding light on the limitations of current approaches to deal with length generalization in algorithmic tasks with transformers.
@@ 2 Related works
Systematic Generalization A consistent stream of research investigates the extent to which neural networks can learn to reason in a systematic and compositional way [11, 30]. Several benchmarks have been proposed to measure this capacity: among the most popular, ListOps [22] and CTL [18] can be seen as instances of formula simplification problems, in which the learning system is tasked to produce the evaluation of a formula which typically requires to compute intermediate results, either implicitly or explicitly. Other formula simplification benchmarks include derivation and integration problems [16] or polynomial simplification [1]. Modern neural architectures achieve different degrees of systematic compositional generalization [7, 15, 23]. Interestingly, recent work has shown the emergence of specialized modules in Mixture of Experts architectures [21], suggesting that modular architectures can outperform monolithic ones in both in- and out-of-distribution testing conditions.
Rewriting systems and neural networks The idea of implementing neural architectures inspired by symbolic rewriting systems was initially explored in unsupervised learning settings, where individual network weights represented tokens to be rewritten [13]. Others implemented a rewriting system for algebraic problems using custom feature engineering methods and a feed-forward network [3], while in [4] the authors propose a reinforcement learning-based system that can learn a general rewriting mechanism by first choosing a region to simplify and then an appropriate rewriting rule.
LLMs Research on prompt engineering techniques has shown that appropriate prompting methods can boost the reasoning capabilities of LLMs [32, 33]. In particular, Chain-of-Though prompting improves LLMs’ accuracy on reasoning tasks by leveraging their auto-regressive nature: the models are elicited to give the final result using multi-step reasoning chains, which enable to process contextual information more effectively. Among the reasoning tasks on which Large Language Models often exhibit impressive problem-solving capabilities we find ‘symbolic reasoning tasks’. Some examples are coin flip, last letter concatenation [33] and boolean variable assignment [2]. These tasks are remarkably similar in spirit to those we consider in this work, since problem instances are synthetically generated and can be solved by applying simple algorithms.
@@ 3 Formula simplification problems
Many systematic generalization benchmarks require to evaluate a possibly very complex formula by replacing expressions with elementary operands that have an equivalent value. We call such problems ‘formula simplification problems’. For example, in ListOps an operation on a list of integers can be replaced with a single integer that is the result of the operation. In order to use formula simplification problems as a benchmark to assess the systematic generalization abilities of neural architectures, we outline a formal framework that includes the class of problems we consider in this work.
@@
In formula simplification problems, we can see formulas $f\in F$ as entities that are composed of two semantically distinct elements: operators $o\in O$ and arguments $a\in A$.  [^1] We also sometimes refer to arguments as operands. We prefer to use the term arguments here to avoid confusion with operators $O$. Arguments can be either atomic elements $e\in E$, such as integers, which are also the final values of any formula, or other formulas: therefore, in general, formulas have a recursive structure (see Figure 1). Furthermore, since in these problems each formula has always one and only one final value, any formula has the generic compact form $f=o(a_{1},...,a_{n})=e$, where $o\in O$, $e\in E$ and $a_{j}\in F\cup E,~{}\forall j\in[1,n]$. Finally, we can define leaf formulas $F^{L}\subset F$ as the subset of formulas whose arguments are all atomic elements: $f^{L}=o(a_{1},...,a_{n})\;\ \mathrm{s.t.}\;\ a_{k}\in E,~{}\forall k\in[1,n]$.
@@
The simplest algorithm that can be applied to solve formula simplification problems is to solve leaf formulas one by one until the formula is completely reduced to an atomic element $e$. Therefore, for any problem, we can define a set of rewriting rules $r\in R$ which map leaf formulas to their values: $r:F^{L}\rightarrow E,~{}\forall r\in R$. Taking into account the problems we consider in this work, we can further characterize the rewriting rules in $R$. First, the rules apply to contiguous sub-sequences of symbols. Second, each rewriting rule is contractive; that is, it maps sub-sequences to shorter ones. Third, the rewriting rules have mutually exclusive applicability scope, i.e. given a leaf formula, one and only one rule can be applied to reduce it. Fourth, several leaf formulas could be present in a formula at once, but the order in which rewriting rules are applied on them is irrelevant.
The notation introduced above allows us to outline a simple algorithm to solve nested formulas, consisting of the iterative application of the composition of three functions: $sel:F\rightarrow F^{L}$, that maps a formula to a leaf formula appearing in it; $sol:F^{L}\rightarrow E$ that solves the leaf formula, i.e. applies an appropriate rewriting rule to map it to an equivalent atomic element; $com:F\times F^{L}\times E\rightarrow F$, that combines the initial formula $f$ and the solution of the leaf formula $f^{L}$ by replacing $f^{L}$ with $sol(f^{L})$ in the initial formula. Notice that the function $sel$ can be applied again on the output of the function $com$ which is a valid formula, thus enabling the implementation of an iterative solution procedure.
Throughout this work, we will use this simple algorithm as a reference to compare the way different models solve formula simplification problems and to critically evaluate their results. Furthermore, we propose a neural implementation of this algorithm in the form of a modular neural architecture called the Neural Rewriting System.
@@ 4 Neural Rewriting System
Figure 1: Examples of solution of ListOps, arithmetic and algebraic formulas. Formulas $f$ are reduced to atomic values $e$ by iteratively solving leaf formulas $f^{L}$, highlighted in yellow.
In this section, we describe the proposed Neural Rewriting System, a neural architecture composed of three integrated modules, called Selector, Solver, and Combiner, each being the neural implementation of one of the functions that can be used to solve formula simplification problems according to the algorithm described in Section 3. The modules are trained independently of each other and interact at test time to solve nested formulas iteratively. The model is designed to achieve strong out-of-distribution generalization, i.e. to solve very deep formulas while being trained on a set of much shallower formulas. We describe the composition of the datasets used in the development of the model in Section 5.2.1. A schematic representation of the system is shown in Figure 2.
4.1 Selector Module
The Selector module implements the function $sel:F\rightarrow F^{L}$, i.e. it is trained to map a formula to a leaf formula appearing therein. In analogy to what happens in humans when they deploy object-based attention to locate algebraic sub-expressions that can be simplified [19], the Selector is trained to identify the last leaf formula [^2] To enable generalization on formulas with more arguments than seen during training, if the identified leaf formula has more than two arguments, we train the Selector to output the smallest fraction of the leaf formula which can be simplified. This corresponds to the operator and the two first arguments appearing in the leaf formula. occurring in the input formula on which a rewriting rule can be applied. When training the Selector, we assume that it will always receive syntactically correct formulas, i.e. we use teacher forcing.
We frame the problem as a sequence-to-sequence task and use a variant of the transformer encoder-decoder [31] to implement the Selector. We modify the vanilla transformer in two ways. First, given the class of problems we consider in this work, we can assume that the problem the Selector needs to solve has a local nature: the leaf formula to be simplified is always located in a limited region of the input sequence. We therefore mask all entries in the self-attention matrix of the encoder but the ones around the main diagonal (i.e., we make them $-\mathrm{inf}$). The diagonal window is $12k+$ 1 tokens wide, where $k$ is a hyperparameter. We also conducted preliminary experiments showing that models with vanilla self-attention achieved worse out-of-distribution generalization. Second, to enable generalization to out-of-distribution problem instances, we require the Selector to identify a leaf formula independently of the length of the input sequence, provided that it is syntactically correct. Recent research on length generalization in transformers provided evidence that their ability to generalize on longer samples can be influenced by the choice of positional encodings, especially when, at test time, these fall out of the range observed during training [5, 29]. We thus use Label-based Positional Encodings [17] to enable the Selector to identify leaf formulas in very long sequences. The positional information of an input sequence of $L$ tokens is thus encoded in the following way: given a sequence of $N$ sinusoidal positional encodings, where $N$ is a large number that represents the maximum expected length of an input, $L$ integers are sampled in the interval $[0,N-1]$ and then sorted. The encodings found in the positions corresponding to the sampled integers are then summed to the embeddings of the tokens in the input sequence.
@@ 4.1 Selector Module
Multi-output generation Label positional encodings introduce randomness in the representation of input sequences. We have observed experimentally that to mitigate the effect of randomness it can be useful to repeat the auto-regressive generation of transformer outputs. After sampling several output sequences from the probability distribution derived from the decoder’s outputs, we choose the best one considering both a measure of confidence of the Selector and a measure of input-output agreement computed by the Combiner.
@@
Given the specialized purpose of the Selector, we can see each output of the module as a candidate leaf formula $\hat{f}^{L}$. We generate any token $\hat{f}_{i}^{L}$ in an output sequence $\hat{f}^{L}$ by sampling from the probability distribution obtained applying the $\mathrm{softmax}$ function to the logits produced by the final fully-connected layer of the decoder. We do not use any temperature parameter when sampling the output tokens. For any input formula $f$, we repeat the stochastic generation process $M$ times, thus generating a sequence of candidate leaf formulas $\hat{F}^{L}=\langle\hat{f}^{L,1},...,\hat{f}^{L,M}\rangle$. We define the confidence of the Selector on any $\hat{f}^{L,j},1\leq j\leq M$ as the joint probability of sampling its tokens: $c(\hat{f}^{L,j})=\prod_{i=1}^{N}p_{i}^{j}$, where $N$ is the number of tokens in $\hat{f}^{L,j}$, and $p_{i}^{j}$ is the probability to sample token $\hat{f}_{i}^{L}$ in $\hat{f}^{L,j}$. We also define an agreement score $a(\hat{f}^{L,j},f)\in[0,1]$ which gives information about the fraction of $\hat{f}^{L,j}$ which is exactly present in the input formula $f$. This measure is computed by the Combiner and thus it is formally defined in Section 4.2. We then select the final output $f^{L}$ of the Selector as the one with the highest Selector confidence which has an agreement score equal to $1$ — that is, it matches the input sequence exactly. More formally, $f^{L}=\hat{f}^{L,j}\in\hat{F}^{L}\;$ $\mathrm{s.t.}\;$ $c(\hat{f}^{L,j})\geq c(\hat{f}^{L,k})~{}\forall j,k\in[1,M]\land a(\hat{f}^{L,j},f)=1$.
@@
Dynamic windowing We implement a dynamic windowing mechanism on longer input sequences that allows us to increase the model’s generalization capacity on complex problem instances. The core idea behind this mechanism is to repeat the process of selecting a leaf formula several times, changing each time the window of the input formula that the Selector observes, and then relying on the confidence $c(\hat{f}^{L})$ to pick the best output. We apply this mechanism on top of multi-output generation by modifying its behavior for sequences longer than a given threshold $T$. Given an input formula $f$, if $|f|<T$ the computation is executed as described before. Otherwise, we generate $M$ copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$, whose lengths will be reduced by applying a window function $w$. Considering any input $f$ as a sequence $f_{1},...,f_{N}$ of $N$ tokens, we define the window function $w(f,k)=f_{k+1},...,f_{N}$ which reduces the length of the input by giving as output its last $k$ tokens. Since the Selector is trained to output the last leaf formula appearing in the input, the window function reduces the input length starting from the first tokens. We divide the sequence of copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$ into 20 groups $F^{(1)},...,F^{(20)}$ of equal size. Intuitively, in each group the length of the input is reduced by a different percentage of tokens. More formally, the window function will be parameterized by $k=\mathrm{floor}(|f^{(i)}|\cdot\frac{j}{20})~{}\forall f^{(i)}\in F^{(j)},~{}\forall j\in[1,20]$. We then pick the final leaf formula using the confidence and agreement scores, as described in the previous paragraph. This ensures that the model can observe the whole input sequence and select a leaf expression in the part of the input where it can identify one with more confidence.